OpenStudy (anonymous):

Find the value of A so that the equation y'-x^6y-x^6=0 has a solution of the form y(x)=A=Be^(x7/7) for any constant B. wat.

OpenStudy (anonymous):

*A+Be^(x^7/7)

ganeshie8 (ganeshie8):

multiply the equation thru by IF : \(\large e^{-\frac{x^7}{7}}\)

ganeshie8 (ganeshie8):

first two terms reduce to form (fg)' which is easily integrable to fg

OpenStudy (anonymous):

thanks. :>

ganeshie8 (ganeshie8):

wat did u get for final solution ha ?

OpenStudy (anonymous):

-1!

ganeshie8 (ganeshie8):

i knw u gave up lol

OpenStudy (anonymous):

uhhh i didnt actually. I didnt get the way you were explaining though so i just did it the original way i was doing and found out i had just made a calculation error the first time around.

ganeshie8 (ganeshie8):

thats fine :) may i knw the final solution u got ?

OpenStudy (anonymous):

I got A = -1.

ganeshie8 (ganeshie8):

Okay, could you shine some light on the "original way" you talking about ? The only way I ever learned about solving these was by multiplying IF (integrating factor) : \( e^{\int p dx} \) It converts y' and y terms to \((fg)'\) form, so that we can use the fundamental theorem calculus : \(\int (fg)' = fg\)

ganeshie8 (ganeshie8):

Here is the solution to present problem using Integrating Factor method : \(\large y'-x^6y-x^6=0 \) multiply the IF \(\large e^{\frac{-x^7}{7}}\) thru out : \(\large e^{\frac{-x^7}{7}}y' -e^{\frac{-x^7}{7}} x^6y - e^{\frac{-x^7}{7}} x^6 = 0\) \(\large e^{\frac{-x^7}{7}}y' -e^{\frac{-x^7}{7}} x^6y = e^{\frac{-x^7}{7}} x^6 = 0\) \(\large (e^{\frac{-x^7}{7}}y)' = e^{\frac{-x^7}{7}} x^6 \) Integrating both sides : \(\large \int (e^{\frac{-x^7}{7}}y)' dx = \int e^{\frac{-x^7}{7}} x^6 dx\) \(\large e^{\frac{-x^7}{7}}y = \int e^{\frac{-x^7}{7}} x^6 dx\) right side is trivial to integrate

ganeshie8 (ganeshie8):

\(\large e^{\frac{-x^7}{7}}y = \int e^{\frac{-x^7}{7}} x^6 dx \) \(\large e^{\frac{-x^7}{7}}y = -e^{\frac{-x^7}{7}} + c \) \(\large y = -1 + ce^{\frac{x^7}{7}} \)

ganeshie8 (ganeshie8):

^^ gives -1 as well, but for some reason i dont see how we can do this in any other way :o

OpenStudy (anonymous):

I just took the derivative of y, subtracted x^6 time y minus x^6 and ended up with this \[Bx^6e^\frac{ x^7 }{ ^7 }-x^6(A+Be ^\frac{ x ^7}{ ^7 })-x^6\] Then i factored out x^6 and was able to cancel Be^x^7/7 with -Be^x^7/7 So i was just left with -A-1=0 -A=1 A=-1

ganeshie8 (ganeshie8):

Awesome ! I see u worked it backwards ! thats a cool trick :)

OpenStudy (anonymous):

woooo! :) thanks for explaining it an additional way.

ganeshie8 (ganeshie8):

:) i see a third simpler way lol

ganeshie8 (ganeshie8):

but i guess wat you did is the simplest way possible !

ganeshie8 (ganeshie8):

I'll put the third method anyways :-

ganeshie8 (ganeshie8):

Method 3 :- \(y'-x^6y-x^6=0 \) \(y' -x^6(y-1) = 0\) \(y' = x^6(y-1)\) \(\large \frac{dy}{dx} = x^6(y-1)\) \(\large \frac{dy}{y-1} = x^6dx\) Integrating both sides : \(\large \int \frac{dy}{y-1} = \int x^6dx\) \(\large \ln (y-1) = \frac{x^7}{7} + c\) \(\large y-1 = Be^{\frac{x^7}{7}}\)

ganeshie8 (ganeshie8):

this method may feel easy to work out..

ganeshie8 (ganeshie8):

*typo it should be \((y+1)\) all the way